OpenStudy (anonymous):

The product of 9 and a number is 12 less than three times that number. What is the number?

OpenStudy (sleepyjess):

Lets make this into an equation, have any ideas?

OpenStudy (sleepyjess):

Ok, so we have the product of 9 and a number, 9x. Then, is 12 less, -12. Than three times that number, = 3x. So overall we get 9x-12=3x. Does that make sense?

OpenStudy (sleepyjess):

Did that make sense, it is ok if no.

OpenStudy (anonymous):

i didn't understand that sorry

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

The product of 9 and a number. That's multiplying 9 with an known number, which we can call 'x'. So we can write that as '9x'. The 'is' separates the two halves of the question with an equal sign(=). So far we have 9x = '12 less than three times that number' 12 less than means that we will subtract 12 from whatever comes next. three times that number, which is the same thing as 3x, so we subtract 12 from '3x', giving us 3x - 12. So our full equation is: 9x = 3x - 12 Can you solve that for 'x'?

OpenStudy (igreen):

Actually I think I did something wrong..lol

OpenStudy (igreen):

Yeah..I think what you did was correct @sleepyjess

OpenStudy (sleepyjess):

Originally I thought the same thing @iGreen

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

@sleepyjess @iGreen

OpenStudy (anonymous):

is this the correct answer @iGreen or @sleepyjess ?

OpenStudy (sleepyjess):

What did you do to get -3?

OpenStudy (anonymous):

sleepyjess please answer i need to know if i am correct

OpenStudy (anonymous):

i did 9x-12=-3x

OpenStudy (sleepyjess):

Why -3x?

OpenStudy (sleepyjess):

I gave you the equation, 9x-12=3x

OpenStudy (sleepyjess):

You have to solve for x

OpenStudy (anonymous):

I GOT IT x=2

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

Yep!

OpenStudy (igreen):

Yep, you got it.. x = 2.

OpenStudy (igreen):

9x - 12 = 3x Add 12 to both sides: 9x = 3x + 12 Subtract 3x to both sides: 6x = 12 Divide 6 to both sides: x = 2

OpenStudy (igreen):

Great work. @sleepyjess

OpenStudy (sleepyjess):

@alpha_female02 , will you please close this question?

OpenStudy (sleepyjess):

Also, \({\huge\bf\color{red}\bigstar\color{orange}\bigstar\color{goldenrod}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{darkgreen}\bigstar\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar\bigstar\bigstar}\color{#11c520}{\bigstar\bigstar\bigstar}}\\{\huge\bf\color{red}W\color{orange}E\color{goldenrod}{L}\color{yellow}{C}\color{lightgreen}{O}\color{darkgreen}M\color{turquoise}{E~}\color{royalblue}{T}\color{purple}{O}~~\color{#00bfff}{Open}\color{#11c520}{Study}}\\{\huge\bf\color{red}\bigstar\color{orange}\bigstar\color{goldenrod}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{darkgreen}\bigstar\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar\bigstar\bigstar}\color{#11c520}{\bigstar\bigstar\bigstar}}\\\small\cal Made~by~@sleepyjess\)